OpenStudy (anonymous):

Derivative of 5/(2x)^3 +2cosx. Step by step please

OpenStudy (anonymous):

second term you do by knowing that the derivative of \[\cos(x)\] is \[-\sin(x)\] so the second part will be \[-2\sin(x)\]

OpenStudy (anonymous):

first part we can do by the "power rule" after rewriting \[\frac{5}{(2x)^3}=\frac{5}{8x^3}=\frac{5}{8}x^{-3}\] the derivative will then be \[-3\times \frac{5}{8}x^{-4}=\frac{-15}{8x^4}\]

OpenStudy (anonymous):

Thanks a ton

OpenStudy (anonymous):

yw